1. Calculate the money market yield and bond equivalent yield of a $100,000 30-day US Treasury bill that is trading at a discount of 5%.

A. The money market yield is 5% and the bond equivalent yield is lower than the money market yield.

B. The money market yield is 5.02% and the bond equivalent yield is higher than the money market yield.

C. The money market yield is 5.09% and the bond equivalent yield is lower than the money market yield.

2. I: Staggered voting: directors are divided into classes in order to limit the ability of shareholders to replace a full board quickly.

II: Majority voting: is required for ordinary resolutions such as appointing an auditor.

III: Supermajority voting: is where more than 50% of the vote is needed, required for important events such as mergers and acquisitions.

A. Statement I is incorrect, Statement II and III are correct.

B. Statement I is correct, Statement II and III are incorrect.

C. All statements are incorrect.
Answer #1

Money market yield = [(Face value – Purchase price)/Purchase price] x (360/Time to maturity)

Face Value = 100,000

Discount = 5% for one year or 360 days hence for 30 days it is 5%*30/360 = 0.4167%

Discount = 0.4167% * 100,000

= 416.7

Purchase Price = 100,000 - 416.7 = 99583.3

Money Market Yield = (100,000 - 99583.3)/ 99583.3 *[360/30]

= 416.7 / 99583.3 * 12 = 5.02%

Bond Equivalent yield = [(Face value – Purchase price)/Purchase price] x (360/Time to maturity)

= (100,000 - 99583.3)/ 99583.3 *[365/30]'

= 416.7 / 99583.3 * 12.167 = 5.09% ( all calculations done above)

Answer ) = Option B , The money market yield is 5.02% and the bond equivalent yield is higher than the money market yield.

answer 2)

option B